# Transport Properties of QGP within a Bayesian Holographic QCD Model

### v1

@@ Abstract
Using a holographic QCD model augmented by Bayesian inference, we calculate key transport coefficients of the quark-gluon plasma (QGP)—including the drag force, jet quenching parameter, heavy quark diffusion coefficient, and shear and bulk viscosities—at finite temperature and chemical potential. Posterior parameter distributions at the 68% and 95% confidence levels (CL), as well as the maximum a posteriori (MAP) estimates, are employed to quantify uncertainties. Our findings indicate that the diffusion coefficient within the Bayesian credible regions aligns with lattice QCD results for $T\sim 1.2T_{c}$ to $2T_{c}$, and is consistent with ALICE experimental measurements near $T_{c}$. The jet quenching parameter obtained from the Bayesian analysis agrees with RHIC and LHC data, while viscosity coefficients show compatibility with existing literature. These results demonstrate the efficacy of a Bayesian holographic approach in elucidating the nonperturbative transport properties of QCD matter.
@@ I Introduction
In this study, we extend a holographic QCD model with Bayesian inference to compute the drag force, jet quenching parameter, heavy quark diffusion coefficient, and viscosities at finite temperature and chemical potential. Employing posterior parameter distributions at 68%, 95% confidence levels (CL) and maximum a posteriori (MAP) estimates, we compare our results with lattice QCD, experimental data, and perturbative predictions, thereby providing a comprehensive analysis of the transport properties of the QGP [111, 110, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132].
@@ II Holographic Model via Einstein-Maxwell-Dilaton Gravity
$A(z)$ is set to mimic the correct behavior of entropy and constrain the temperature-dependent model. The function $f(z)$ describes the dependence of model on the chemical potential, which is fixed by the baryon number susceptibility. Then, we can derive
@@
There are three undetermined parameters $a$, $b$, and $d$ in $A(z)$, and two parameters $c$ and $k$ in $f(z)$, along with the Newton constant $G_{5}$, making in total a six-dimensional parameter space. The six parameters, as determined by Bayesian inference from the data of the equation of state and baryon number susceptibility from the lattice QCD, for the 2+1 flavor system, we used Bayesian inference to determine the range of values and the MAP values for these six parameters, as shown in Table 1 [171].
@@
Table 1: In the 2+1 flavor system, the 95% CL ranges and MAP values for the six model parameters aa\;,bb\;,cc\;,dd\;,kk\; and $G_{5}$ are determined through Bayesian inference [171].
@@ III Heavy Quark Drag Force
where $\sqrt{\lambda}=\frac{g_{YM}^{2}N_{c}}{4\pi}=\frac{L^{2}}{\alpha^{\prime}}$ is the ’t Hooft coupling. The only one new parameter $\alpha^{\prime}$ is taken to be 2 in this paper. $z_{s}$ is the radial location of the string world sheet horizon [111], which is obtained as the numerical solution of the following equation [175]
@@
According to Eq. (12), it can be derived that energy loss is equal to the drag force, which allows us to plot the relationship between energy loss and momentum in Bayesian inference. Fig. 3 illustrates the relationship between the bottom quark ($m_{b}=4.7\enspace\rm GeV$) and charm quark ($m_{c}=1.3\enspace\rm GeV$) in a 2+1 flavor system at zero chemical potential and temperature $T=T_{c}$ [38, 39]. Fig. 4 shows the relationship between the bottom quark and charm quark in a 2+1 flavor system at zero chemical potential and temperature $T=2T_{c}$, where the light-colored area represents the range of energy loss under 68% CL and 95% CL, and the curve line indicates the energy loss value at the MAP values. Fig. 3 (a) and Fig. 3 (b) represent the 95% CL and 68% CL, respectively, and the same applies to Fig. 4. From Fig. 3 and Fig. 4, it is evident that energy loss increases with momentum. The mass of the quarks also affects energy loss; lighter quarks result in greater energy loss. Additionally, a comparison between Fig. 3 and Fig. 4 shows that temperature has a more significant impact on energy loss than quark mass, with higher temperatures leading to increased energy loss.
Figure 3: Energy loss of the bottom quark ($m_{b}=4.7\enspace\rm GeV$) and charm quark ($m_{c}=1.3\enspace\rm GeV$) as a function of momentum $p$ (in GeV) at temperature $T=T_{c}$. (a) The gray area represents the relationship between energy loss and momentum at the 95% CL. (b) The gray area represents the relationship between energy loss and momentum at the 68% CL. The red and blue curve shows the results at the MAP values.
Figure 4: Energy loss of the bottom quark ($m_{b}=4.7\enspace\rm GeV$) and charm quark ($m_{c}=1.3\enspace\rm GeV$) as a function of momentum $p$ (in GeV) at temperature $T=2T_{c}$. (a) The gray area represents the relationship between energy loss and momentum at the 95% CL. (b) The gray area represents the relationship between energy loss and momentum at the 68% CL. The red and blue curve shows the results at the MAP values.
@@ IV Heavy Quark Diffusion Coefficient
We compared the spatial heavy quark diffusion coefficient, normalized by $2\pi T$, with estimates from lattice QCD, ALICE experiments, and Next-to-Leading Order (NLO) perturbative predictions, as depicted in Fig. 6, it can be seen that our model’s 68% CL, 95% CL and MAP values results almost fall within the error bars of the lattice data for $N_{f}=2+1$ [177]. Our results are also in very good agreement with the results from the ALICE experiment [178]. At high temperatures, our values of $2\pi TD$ are basically consistent with the NLO perturbative predictions [179]. Additionally, our Bayesian inference completely coincide with the region of the Duke hydro/transport model [180].
Figure 6: The scaled diffusion coefficient $2\pi TD$ as a function of $T/T_{c}$ , with the quark velocity set to $v=0.3$. The lattice data in the figure are results from Ref. [177]. The results of ALICE are from Ref. [178]. The orange dashed line represents the NLO perturbative calculations [179]. The pink region represents Bayesian inference results from the Duke hydro/transport model [180]. The red curves and gray area represents the results of our model. (a) The gray area represents the variation of $2\pi TD$ with $T/T_{c}$ at the 95% CL; (b) the gray area represents the variation of $2\pi TD$ with $T/T_{c}$ at the 68% CL. The red curve shows the variation of $2\pi TD$ with temperature at the MAP values.
@@ VI Bulk viscosities
$\zeta=\lim_{\omega\to 0}\frac{1}{\omega}\text{Im}\,G^{R}_{T^{i}_{i},T^{j}_{j}}(\omega,\mathbf{0}),$
where $G^{R}_{T^{i}_{i},T^{j}_{j}}$ is the retarded Green’s function of the trace of the spatial stress-energy tensor. According to the holographic dictionary, the Green’s function of the stress tensor can be extracted through metric perturbations (e.g., $g_{\mu\nu}\to g_{\mu\nu}+h_{\mu\nu}$). When applying the Kubo formula, we choose the spatial components of the momentum as $\vec{q}=0$ and assume that $h_{\mu\nu}$ depends only on time $t$ and the $z$-direction, i.e., $h_{\mu\nu}=h_{\mu\nu}(t,z)$. These metric perturbation components should include $h_{xx}$, $h_{yy}$, and $h_{zz}$. We assume spatial rotational symmetry, so $h_{xx}=h_{yy}=h_{zz}$. Following this reasoning, the bulk viscosity can be extracted from the retarded Green’s function of the scalar mode. The imaginary part of the retarded Green’s function $G^{R}(\omega)$ is related to the conserved flux $\mathcal{F}(\omega)$ via
$\text{Im}\,G^{R}(\omega)=-\frac{\mathcal{F}(\omega)}{4\pi G_{5}},.$
@@
$\mathcal{F}(\omega)=\frac{L^{3}e^{3A(z)}g(z)}{4z^{2}(A^{\prime}-1/z)^{2}}\left|\text{Im}\left(h_{xx}^{*}h_{xx}^{\prime}\right)\right|.$
A notable feature of the perturbation equations of motion is that the equation for the $h_{xx}$ component decouples completely from $h_{zz}$ and $h_{tt}$. The equation of motion for the perturbation $h_{xx}$ takes the form:
@@
The temperature dependence of the entropy density and bulk viscosity is shown in Figs. 11 (a) and (b), respectively. In Figs 11 (a) and (b), the gray regions correspond to the 95% CL, the red curves represent the temperature dependence of the entropy density and bulk viscosity calculated using the MAP values. It can be observed that the model aligns well with the 90% CL of the JETSCAPE Bayesian model [182] at higher temperatures, as well as the findings from the Duke University team [168].
@@ VII Shear viscosity with higher derivative corrections
Here, $c_{0}=\frac{L^{2}e^{2A_{h}}}{z_{h}^{2}}\frac{1}{2g^{\prime}A^{\prime}}=\frac{L^{2}e^{2A_{h}}}{z_{h}^{2}}\frac{1}{2g^{\prime}(A^{\prime}-\frac{1}{z_{h}})}$. When $\beta=0.01$ and $\gamma=-\sqrt{1/2.9}$, Fig. 12 shows the functional dependence of the $\eta/s$ ratio on $T/T_{c}$. It can be observed that the $\eta/s$ ratio exhibits a dip near $T=1.2T_{c}$. The temperature corresponding to this minimum value almost coincides with the temperature corresponding to the peak of the bulk viscosity. In Fig. 12, the blue band represents the 95% CL, while the black curve shows the functional dependence of the $\eta/s$ ratio on $T/T_{c}$ at the MAP value. Fig. 12 (a) displays a magnified view of the results shown in Fig. 12 (b), where our findings are compared with various model predictions [183, 14, 184, 19, 182, 185, 186].
@@ VIII Summary
By integrating a holographic Einstein-Maxwell-dilaton model with Bayesian inference, we have quantitatively analyzed the transport properties of the QGP at finite temperature and chemical potential. Our results for the drag force, diffusion coefficient, jet quenching parameter, and viscosities are qualitatively consistent with Lattice QCD, experimental measurements, and perturbative predictions within uncertainties. This work demonstrates the power of Bayesian calibration of holographic models to extract nonperturbative QCD dynamics reliably. Future work could extend this framework to include finite baryon density effects, anisotropic backgrounds, or higher-derivative corrections to further refine the understanding of strongly coupled QCD matter.

### v2

@@ Abstract
Using a holographic QCD model augmented by Bayesian inference, we calculate key transport coefficients of the quark-gluon plasma (QGP)including the drag force, jet quenching parameter, heavy quark diffusion coefficient, and shear and bulk viscositiesat finite temperature and chemical potential. Posterior parameter distributions at the 68% and 95% confidence levels (CL), as well as the maximum a posteriori (MAP) estimates, are employed to quantify uncertainties. Our findings indicate that the diffusion coefficient within the Bayesian credible regions aligns with lattice QCD results for $T\sim 1.2T_{c}$ to $2T_{c}$, and is consistent with ALICE experimental measurements near $T_{c}$. The jet quenching parameter obtained from the Bayesian analysis agrees with RHIC and LHC data, while viscosity coefficients show compatibility with existing literature. These results demonstrate the efficacy of a Bayesian holographic approach in elucidating the nonperturbative transport properties of QCD matter.
@@ I Introduction
In this study, we extend a holographic QCD model with Bayesian inference to compute the drag force, jet quenching parameter, heavy quark diffusion coefficient, and viscosities at finite temperature and chemical potential. Employing posterior parameter distributions at 68%, 95% confidence levels (CL) and maximum a posteriori (MAP) estimates, we compare our results with lattice QCD, experimental data, and perturbative predictions, thereby providing a comprehensive analysis of the transport properties of the QGP [111, 110, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133].
@@ II Holographic Model via Einstein-Maxwell-Dilaton Gravity
$A(z)$ is set to mimic the correct behavior of entropy and constrain the temperature-dependent model. The function $f(z)$ describes the dependence of model on the chemical potential, which is fixed by the baryon number susceptibility. The constant k corresponds to an overall constant rescaling of the Maxwell term near the conformal boundary. Such a constant prefactor does not change the gauge invariant; it can be absorbed by a field redefinition of the bulk gauge potential. Then, we can derive
@@
There are three undetermined parameters $a$, $b$, and $d$ in $A(z)$, and two parameters $c$ and $k$ in $f(z)$, along with the Newton constant $G_{5}$, making in total a six-dimensional parameter space.
The six parameters $\theta=(a,b,c,d,k,G_{5})$ used in our model are extracted from Ref. [172]. That work employs a Bayesian analysis within the EMD framework, constrained by lattice QCD data for the thermodynamic quantities $(S/T^{3},\chi_{2}^{B},C_{s}^{2})$ at zero chemical potential [178, 179], to determine the posterior distribution of the model parameters. The specific analytical procedure is as follows: first, the parameter inference framework is established based on Bayes’ theorem:
$\mathit{P}(\boldsymbol{\theta}|\mathrm{data})\propto\mathit{P}(\mathrm{data}|\boldsymbol{\theta})\mathit{P}(\boldsymbol{\theta})$ (12)
Within the Bayesian framework, the posterior distribution $\mathit{P}(\theta|\mathrm{data})$ describes the probability distribution of the model parameters $\theta$ after incorporating the observed data. The prior distribution $\mathit{P}(\theta)$ represents the initial belief or existing knowledge about $\theta$ before observing the data. The likelihood function $\mathit{P}(\mathrm{data}|\theta)$ quantifies the plausibility of observing the current data given specific values of $\theta$. Here, ”data” specifically refers to lattice QCD data, including $S/T^{3}$, $\chi_{2}^{B}$, and $C_{s}^{2}$, while $\theta$ corresponds to the six parameters in the EMD model.
Regarding the prior distribution $\mathit{P}(\theta)$, the parameter set $\theta=(a,b,c,d,k,G_{5})$ for the EMD model was assigned prior ranges based on existing research [69], as detailed in Table I of Ref. [172]. Following these ranges, we generated 300 parameter samples using the Latin Hypercube Sampling(LHS) [180, 181] method. These parameters were then substituted into the EMD model to compute the corresponding values of $S/T^{3}$, $\chi_{2}^{B}$, and $C_{s}^{2}$. Subsequently, the results were subjected to dimensionality reduction via Principal Component Analysis (PCA) [182]. Finally, a Gaussian Process emulator [183] was constructed based on the 300 parameter sets and their corresponding PCA-reduced outputs. This emulator will be used to establish the likelihood function in subsequent steps.
Prior
Parameter min max
$a$ 0.110 0.310
$b$ 0.005 0.031
$c$ -0.280 -0.205
$d$ -0.240 -0.110
$k$ -0.910 -0.770
$G_{5}$ 0.375 0.430
Table 1: The prior parameter ranges for the EMD model [172].
Furthermore, for the likelihood function $P(\mathrm{data}\mid\theta)$, we choose to adopt a Gaussian distribution:
$\mathit{P}(\mathrm{\mathrm{data}}|\boldsymbol{\theta})=\underset{i}{\prod}\frac{1}{\sqrt{2\pi}\sigma_{i}}e^{-\frac{[y_{i}(\boldsymbol{\theta})-y_{i}^{\text{lattice}}]^{2}}{2\sigma_{i}^{2}}}$ (13)
The term $y_{i}^{\text{lattice}}$ represents the lattice QCD results, while $y_{i}(\boldsymbol{\theta})$ is the prediction from the EMD model, which is approximated by a Gaussian emulator to enhance computational efficiency. Here, $\sigma_{i}$ encompasses the uncertainties from both the lattice QCD data and the Gaussian emulator. Finally, by sampling the posterior distribution via the Markov Chain Monte Carlo (MCMC) [184, 185], we obtained the posterior distributions for the EMD model parameters. The corresponding 95% Confidence Levels (CL) and the Maximum a Posteriori (MAP) estimates are presented in Table II[172].
@@
Table 2: In the 2+1 flavor QCD system, the distributions of the EMD model parameters $(a,b,c,d,k,G_{5})$ after Bayesian inference, including the maximum a posteriori values and the 95% CL ranges[172].
@@ III Heavy Quark Drag Force
where $\sqrt{\lambda}=\frac{g_{YM}^{2}N_{c}}{4\pi}=\frac{L^{2}}{\alpha^{\prime}}$ is the ’t Hooft coupling. The only new parameter, $\alpha^{\prime}$, is set to 2. Its value is determined by fitting to the NLO perturbative results in the high-temperature region and by minimizing the corresponding deviation. $z_{s}$ is the radial location of the string world sheet horizon [111], which is obtained as the numerical solution of the following equation [176]
@@
According to Eq. (14), it can be derived that energy loss is equal to the drag force, which allows us to plot the relationship between energy loss and momentum in Bayesian inference. Fig. 3 illustrates the relationship between the bottom quark ($m_{b}=4.7\kern 5.0pt\rm GeV$) and charm quark ($m_{c}=1.3\kern 5.0pt\rm GeV$) in a 2+1 flavor system at zero chemical potential and temperature $T=T_{c}$ [38, 39]. Fig. 4 shows the relationship between the bottom quark and charm quark in a 2+1 flavor system at zero chemical potential and temperature $T=2T_{c}$, where the light-colored area represents the range of energy loss under 68% CL and 95% CL, and the curve line indicates the energy loss value at the MAP values. Fig. 3 (a) and Fig. 3 (b) represent the 95% CL and 68% CL, respectively, and the same applies to Fig. 4. From Fig. 3 and Fig. 4, it is evident that energy loss increases with momentum. The mass of the quarks also affects energy loss; lighter quarks result in greater energy loss. Additionally, a comparison between Fig. 3 and Fig. 4 shows that temperature has a more significant impact on energy loss than quark mass, with higher temperatures leading to increased energy loss.
Figure 3: Energy loss of the bottom quark ($m_{b}=4.7\kern 5.0pt\rm GeV$) and charm quark ($m_{c}=1.3\kern 5.0pt\rm GeV$) as a function of momentum $p$ (in GeV) at temperature $T=T_{c}$. (a) The gray area represents the relationship between energy loss and momentum at the 95% CL. (b) The gray area represents the relationship between energy loss and momentum at the 68% CL. The red and blue curve shows the results at the MAP values.
Figure 4: Energy loss of the bottom quark ($m_{b}=4.7\kern 5.0pt\rm GeV$) and charm quark ($m_{c}=1.3\kern 5.0pt\rm GeV$) as a function of momentum $p$ (in GeV) at temperature $T=2T_{c}$. (a) The gray area represents the relationship between energy loss and momentum at the 95% CL. (b) The gray area represents the relationship between energy loss and momentum at the 68% CL. The red and blue curve shows the results at the MAP values.
@@ IV Heavy Quark Diffusion Coefficient
We compared the spatial heavy quark diffusion coefficient, normalized by $2\pi T$, with estimates from lattice QCD, ALICE experiments, and Next-to-Leading Order (NLO) perturbative predictions, as depicted in Fig. 6, it can be seen that our model’s 68% CL, 95% CL and MAP values results almost fall within the error bars of the lattice data for $N_{f}=2+1$ [186]. The gray uncertainty band represents the computed results using the posterior parameters at the 68%CL and 95%CL, while the red curve corresponds to the result obtained from the MAP parameter set. The uncertainty in the gray band originates from the range of the posterior parameters. Our results are also in very good agreement with the results from the ALICE experiment [187]. At high temperatures, our values of $2\pi TD$ are basically consistent with the NLO perturbative predictions [188]. Additionally, our Bayesian inference completely coincide with the region of the Duke hydro/transport model [189].
Figure 6: The scaled diffusion coefficient $2\pi TD$ as a function of $T/T_{c}$ , with the quark velocity set to $v=0.3$. The lattice data in the figure are results from Ref. [186]. The results of ALICE are from Ref. [187].The orange dashed line represents the next-order perturbative calculation with the renormalization scale set to $\mu_{R}=4\pi T$, while the green dashed line represents the next-order perturbative calculation with the renormalization scale set to $\mu_{R}=2\pi T$ [188]. The pink region represents Bayesian inference results from the Duke hydro/transport model [189]. The red curves and gray area represents the results of our model. (a) The gray area represents the variation of $2\pi TD$ with $T/T_{c}$ at the 95% CL; (b) the gray area represents the variation of $2\pi TD$ with $T/T_{c}$ at the 68% CL. The red curve shows the variation of $2\pi TD$ with temperature at the MAP values.
@@ V Jet Quenching Parameter
The parameter $\lambda$(or $\alpha^{\prime}$) is the only new parameter in the paper. At the end of this section, we have added Fig. 11 to illustrate two $\lambda$-sensitive quantities, namely the jet quenching parameter $\hat{q}$ and the diffusion coefficient. As shown in the figure, increasing the value of $\lambda$ leads to an enhancement of $\hat{q}$, while the diffusion coefficient decreases as $\lambda$ increases.
Figure 11: The figure shows two $\lambda$-sensitive quantities: the jet quenching parameter $\hat{q}$ (a) and the diffusion coefficient $D$ (b).
@@ VI Bulk viscosities
$\zeta=\lim_{\omega\to 0}\frac{1}{\omega}\text{Im}\,G^{R}_{T^{i}_{i},T^{j}_{j}}(\omega,\mathbf{0}),$ (27)
where $G^{R}_{T^{i}_{i},T^{j}_{j}}$ is the retarded Green’s function of the trace of the spatial stress-energy tensor. According to the holographic dictionary, the Green’s function of the stress tensor can be extracted through metric perturbations (e.g., $g_{\mu\nu}\to g_{\mu\nu}+h_{\mu\nu}$). When applying the Kubo formula, we choose the spatial components of the momentum as $\vec{q}=0$ and assume that $h_{\mu\nu}$ depends only on time $t$ and the $z$-direction, i.e., $h_{\mu\nu}=h_{\mu\nu}(t,z)$. These metric perturbation components should include $h_{xx}$, $h_{yy}$, and $h_{zz}$. We impose spatial rotational symmetry, which allows us to identify the metric perturbations in the three spatial directions as $h_{xx}=h_{yy}=h_{zz}$. Following this reasoning, the bulk viscosity can be extracted from the retarded Green’s function of the scalar mode. The imaginary part of the retarded Green’s function $G^{R}(\omega)$ is related to the conserved flux $\mathcal{F}(\omega)$ via
$\text{Im}\,G^{R}(\omega)=-\frac{\mathcal{F}(\omega)}{4\pi G_{5}}.$ (28)
@@
$\mathcal{F}(\omega)=\frac{L^{3}e^{3A(z)}g(z)}{4z^{2}(A^{\prime}-1/z)^{2}}\left|\text{Im}\left(h_{xx}^{*}h_{xx}^{\prime}\right)\right|.$ (29)
We fix this residual gauge freedom by choosing the dilaton (radial) gauge, i.e. we use the background scalar to parametrize the holographic direction, which is equivalent to setting $\delta\phi=0$. With this gauge fixing the physical scalar mode is entirely captured by the metric fluctuation $h_{ii}$, and its equation of motion decouples from the remaining perturbations. This procedure follows the standard treatment used, e.g., in Ref. [91] for the computation of bulk viscosity. A notable feature of the perturbation equations of motion is that the equation for the $h_{xx}$ component decouples completely from $h_{zz}$ and $h_{tt}$. The equation of motion for the perturbation $h_{xx}$ takes the form:
@@
The temperature dependence of the bulk viscosity to entropy density ratio is shown in Figs. 12 (a) and (b), respectively. The pronounced peak of bulk viscosity around $T_{c}$ could have important consequences for particle spectra and flow observables. In Figs. 12 (a) and (b), the gray band displays results derived from posterior parameters within the 95% CL, while the red curve is obtained from the MAP parameter estimate. The breadth of the gray band directly reflects the uncertainty propagated from the posterior parameter distribution. It can be observed that the model aligns well with the 90% CL of the JETSCAPE Bayesian model [191] at higher temperatures, as well as the findings from the Duke University team [169].
@@ VII Shear viscosity with higher derivative corrections
Here, $c_{0}=\frac{L^{2}e^{2A_{h}}}{z_{h}^{2}}\frac{1}{2g^{\prime}A^{\prime}}=\frac{L^{2}e^{2A_{h}}}{z_{h}^{2}}\frac{1}{2g^{\prime}(A^{\prime}-\frac{1}{z_{h}})}$. When $\beta=0.01$ and $\gamma=-\sqrt{0.3448}$, Fig. 13 shows the functional dependence of the $\eta/s$ ratio on $T/T_{c}$. It can be observed that the $\eta/s$ ratio exhibits a dip near $T=1.2T_{c}$.
The temperature corresponding to this minimum value almost coincides with the temperature corresponding to the peak of the bulk viscosity. Near the transition region, the trace anomaly exhibits a pronounced peak and the speed of sound $c_{s}^{2}$ reaches a minimum, reflecting the strong breaking of conformal symmetry in QCD. These features enhance the interaction strength, which naturally drives $\eta/s$ toward its minimum. This behavior is consistent with general expectations in strongly coupled gauge theories, where $\eta/s$ approaches the KSS bound in the vicinity of the transition.
In Fig. 13, the blue band represents the results computed using posterior parameters at the 95% CL, while the black curve illustrates the functional dependence of the $\eta/s$ ratio on $T/T_{c}$ at the MAP value. The uncertainty associated with the blue band originates from the range of the posterior parameters. Fig. 13 (a) displays a magnified view of the results shown in Fig. 13 (b), where our findings are compared with various model predictions [192, 14, 193, 19, 191, 194, 195].
@@ VIII Summary
In this work, we performed a comprehensive investigation of the transport properties of the quark–gluon plasma using a five-dimensional Einstein–Maxwell–dilaton (EMD) holographic model whose parameters are fully constrained by Bayesian inference. In contrast to previous holographic studies that rely on specifically chosen potentials or phenomenological inputs, our model incorporates the complete posterior distributions inferred from lattice QCD thermodynamics, thereby enabling a systematic and statistically controlled quantification of theoretical uncertainties for all transport observables. This Bayesian holographic framework provides a unified and nonperturbative approach to studying the strongly coupled dynamics of QCD matter.
We calculated several key transport coefficients—namely the heavy-quark drag force, the spatial diffusion coefficient, the jet quenching parameter, and the bulk and shear viscosities—at finite temperature and chemical potential. These observables probe complementary aspects of the QGP and jointly capture the interplay between thermal effects, nonconformality, and strong coupling.
The drag force was shown to increase monotonically with both temperature and quark velocity, consistent with a medium whose color charge density and interaction strength grow with temperature. The Bayesian uncertainty band demonstrates that this trend is robust against variations in the EMD model parameters and exhibits quantitative agreement with expectations from strongly coupled dynamics. Because the worldsheet horizon depends sensitively on the background metric, the drag force serves as a stringent probe of the gravitational dual and directly reflects the nonconformal structure encoded in the metric deformation.
For the diffusion coefficient, we demonstrated that the scaled quantity $D/D_{\mathrm{SYM}}$ approaches unity at high temperature, indicating an emergent conformal behavior of the plasma in the $T\gg T_{c}$ regime. Near the transition region, however, the diffusion coefficient deviates substantially from its conformal value, signaling strong nonperturbative effects. Remarkably, our Bayesian credible intervals align well with lattice QCD calculations, ALICE experimental estimates, and next-to-leading-order perturbative predictions at high temperatures. This agreement demonstrates the predictive power of the Bayesian EMD model.
The jet quenching parameter exhibits a pronounced peak slightly above the phase transition temperature. This behavior mirrors that found in earlier holographic studies but is now obtained with statistically meaningful uncertainty quantification. The peak provides direct evidence of enhanced transverse momentum broadening near $T_{c}$, where the medium’s color interaction strength is maximal. The magnitude and temperature dependence of $\hat{q}$ derived from our model agree well with phenomenological extractions from RHIC and LHC data, as well as with the HT-BW and HT-M Bayesian extractions from the JET Collaboration.
The viscosities further illuminate the role of strong coupling in the transition region. Consistent with the nonconformal equation of state, we find a sharp peak in the bulk viscosity near $T_{c}$, accompanied by a minimum in the shear viscosity. The near coincidence of these extrema indicates that the rapid change in the trace anomaly around the phase transition significantly alters the dissipative properties of the medium. Once higher-derivative corrections are included, $\eta/s$ acquires a temperature dependence with a well-defined dip, consistent with expectations from hydrodynamic modeling and with Bayesian extractions of $\eta/s(T)$ from the JETSCAPE and Duke analyses. These results highlight that the EMD model—despite its relative simplicity—captures the essential dynamics underlying both conformal and nonconformal transport effects.
Taken together, our findings demonstrate that a Bayesian-calibrated holographic EMD model can consistently and simultaneously describe a wide range of transport coefficients across the entire temperature regime relevant to heavy-ion collisions. The ability to reproduce lattice QCD thermodynamics, match experimental constraints, and quantify theoretical uncertainties represents a significant advancement over traditional holographic approaches. This work therefore provides a robust and predictive nonperturbative framework for studying QCD matter and offers valuable insights into the microscopic mechanisms governing jet energy loss, heavy-quark dynamics, and viscous transport in the strongly coupled quark–gluon plasma. In the future, our results can be used as inputs to transport models, such as the Linear Boltzmann Transport model, to compute observables like $R_{AA}$, $v_{2}$, and others [197, 198, 199].
Acknowledgement